OpenStudy (anonymous):

Evaluate the limit Lim ((2−x)(3+10x))/((3−3x)(4+4x)) x->INF

OpenStudy (anonymous):

Lim = 5/6

OpenStudy (anonymous):

can u explain how ?

OpenStudy (anonymous):

yup, give me a sec

OpenStudy (anonymous):

(2-x)(3+10x)=something -10x2 (3-3x)(4+4x)=something -12x2 In the limit you don't really care about the 'something' since it will grow slower than x2 So Lim = Lim -10x2/-12x2 = -10/-12=5/6

OpenStudy (anonymous):

Or if you really want the whole thing, Lim = Lim (-10x2+17x+6)/(-12x2+12) = Lim -10x2/-12x2 = -10/-12=5/6

OpenStudy (anonymous):

Cheers ;)

OpenStudy (turingtest):

to clarify: Syderic distributed, then divided top an bottom by x^2 ax x -> infty only the constants will remain

OpenStudy (turingtest):

Syderitic, excuse me :P

OpenStudy (turingtest):

*as x-> infty

OpenStudy (anonymous):

no prob ;)

OpenStudy (anonymous):

Thank u all

OpenStudy (anonymous):

one thing about Lim (-10x2+17x+6)/(-12x2+12) how did you solve it?

OpenStudy (anonymous):

hum... well u see: when calculating Lim of p1(x)/p2(x) with p1 and p2 being polynomials you want to see wich values will be the most 'influents' on limit. ex (see image) as can see x3 will be far greater than x2 or x in inf soo we only look at the terms with greater order and ignore the others, since in inf they will be too small when comparing to those of higher order.

OpenStudy (anonymous):

so i ignore (17x+6) and (12) ?